Journalist Report – January 02nd


 Journalist Report
Ben Durkee, Crew 218 Journalist

Sol 11

We're gonna be famous!

We had an extra sugary breakfast of cereal and cookies this morning while we waited for the film crew's spacecraft to arrive. The overload of carbohydrates in our bodies may have made us a little antsy as we marked time in the Hab. Our food supplies are beginning to run on empty, but we still have a surplus of flour and sugar, so Pat and I woke up at 6 AM Mars time to make our confectioneries. Hey, calories are calories. I'll get diabetes on Mars before I starve on Mars.

Our Earthling guests were a tad late - the entry, descent, and landing phase was a little rockier than they expected, I reckon. We in the Hab were worried about our potential acting careers falling through, but it was actually a blessing in disguise as it afforded us more time to prepare for our first EVA of the day. We got everything in order for Pat, Shefali, and Jonathan's excursion, and then our otherworldly visitors knocked at the airlock.

We discussed their desires and our own related to the filming, the most important thing being that the mission is priority one, cinematics are in the passenger seat. We hammered out both party's expectations for the day, and then the crew accompanied our three EVA musketeers to their nearby destination. Their EVA site was just within visual range of the Hab, and it was very entertaining to watch our astronauts walk up and down the same hill for the cameras. They got some footage of Jonathan's meteorological surveys and Pat's seismic tests, and then promptly returned to the Hab.

Since they consumed very little rover charge on their excursion, there was a quick turnaround between EVAs. Those three tagged in the other half of the crew - namely Cesare, LuzMa, and yours truly - and we set off to our new, considerably farther, goal. We had a good spot in mind that was both useful for data collection and photogenic for the cameras. It was the perfect spot, the only issue was the terrain being too harsh for the camera crew's Chevy Spaceship. They had to disembark their craft nearby and walk to the site.

Once we all convened at the picturesque bluff northeast of the Hab, we began analyzing and they began filming. I got some solid radiological data and some fun pictures of the filming crew in action. They were very appreciative of the photos, as they hardly get any shots of themselves behind the scenes. Photographers/videographers are this world's unsung heroes, that's for sure. By far the best thing to come of this EVA, though, is the drone footage of me wiping out in mud while carefully descending the hill. Don't tell the crew (because I am a broke college student), but I would pay good money for that video.

We returned to our rovers and they to their spacecraft. I was very impressed with how they were able to hold their breaths for the entire EVA. We complimented their lung capacity and proceeded back to the Hab for an end-of-day recap and very late lunch. Stories were shared between crews and before we knew it, the time had come for their orbital window back to Earth. We saw them off and lamented how we will be departing the same way soon.

I've started to get sentimental about this place. We've only been here sub-two weeks, but Mars has a way of sucking you in and making you wanna stay. And I'm not just talking about gravity. This mission seems like it has gone by quicker than it took for our dried mango supply to perish. That is to say: very fast.

I have a powerful cognitive weapon at my disposal, though. Any time I start reminiscing about the elements that make this place so grand, I just have to think about the beans. Oh, the notorious beans. We had beans and jambalaya for dinner last night and it was a delicious meal - we also cooked a large surplus of beans for future meals between now and our departure. All in all, a great time. The problems arose when I realized I was on dish duty that night. I looked at the culinary carnage around the table and my stomach dropped - both from the beans eaten and the beans uneaten.

See, beans have an uncanny ability to adhere to dishes. Not only that, but they spread to any dish they even remotely come in contact with. If you take some bean water (yum) and transfer it to other dishes, no matter how heavily you dilute it, the water will somehow always potently remember its bean history. It's like it has Post Traumatic Bean Disorder. It's a nightmare for dishwashing, especially when you are trying to conserve water to the best of your ability. Soon enough our whole kitchen reeked of bean. It propogates like the Bubeanic Plague. I spent multiple hours cleaning those dishes, and still had to call in backup to cleanse them of their blight. It was unbeanlievable. Not to mention our fridge full of leftover beans. Opening that thing is opening Beandora's Box. Don't do it.

I realize I have spent upwards of two paragraphs discussing beans and bean related anguish, but this is the only outlet myself and the rest of the crew have for our deep-seeded pain.

Tomorrow, we have one more EVA before beginning to wrap up.

Tonight, I count beans instead of sheep.

But, you know, other than that our time here has bean phenomenal.

GreenHab Report – January 02nd



Crew 218 GreenHab Report 02-JAN-20

Crew GreenHab Officer: Dr. Jonathan R. Buzan and HSO Shefali Rana

Environmental control: Heating.

Shade cloth (40% and 30%) on.

Average temperature: 21.2°C; 20%


Floor Unit: 15°C

Electronic: 21.1°C

humidity 19%


Floor Unit: 16°C

Electronic: 21.1°C

humidity 19%


Floor Unit: 15°C

Electronic: 20.9°C

humidity 20%

Max: 25.6°C; 22%

Min: 16.8°C; 19%

Hours of supplemental light: Light system 6:00PM-11:00AM.

Daily water usage for crops: 4.3Gal.

Daily water usage for research and/or other purposes: N/A

Water in Blue Tank – ~80 Gal.

Time(s) of watering for crops:

9:40AM: 1.1 Gal. Cucumber and snow peas

4:15PM: 2.1 Gal. Watering the herb bed for herb garden and hanging plants.

6:15PM: 1.1 Gal. Final watering.


Change to crops: None.


1st sprouts: N/A

—Watered cucumber due to dry soil.

—Film crew captured GreenHab activities of Jonathan and Shefali.

—Gnat spray has arrived.

Harvest: N/A

Commander Report – January 02nd



Crew 218 Commander Report 02-01-2020

Sol 11 – Light, camera, action

Today we had a special day: a visit from an Earthling filming crew who spent some time grabbing our activities as analog astronauts. Our day proceeded as usual, with two EVAs after breakfast and a good amount of research. Despite the clouds and the cold that still persist, we had successful excursions, after which our guests had more taste of our research activities and of our social life. Tomorrow we will already be cleaning and preparing the hab for the following crew. As usual, time here flies and it is hard to realize that we actually are in a new year!

Cesare Guariniello, Commander

EVA Report – January 02nd


 EVA #: 10

Author: Pat Pesa

Purpose of EVA: Seismometric measurements; EVA and EMU evaluation project; Filming crew

Start time: 11:17

End time: 12:35

Narrative: Jonathon, Pat, and Shefali took the short Rover trip to the Marble Ritual. They were met there by a guest filming crew to observe research activity. Seismic measurements were taken at the site at several distance intervals.

Destination: Marble Ritual

Coordinates (use UTM NAD27 CONUS): 12S 518800, 4250700

EVA Participants: Pat Pesa, GEO; Jonathan Buzan, SCI; Shefali Rana, HSO

Road(s) and routes per MDRS Map: Cow Dung Road

Mode of travel: Driving and walking

Vehicles you will be using (If applicable): Curiosity and Spirit

EVA #: 11

Author: Cesare Guariniello

Purpose of EVA: weather observation; Geology samples collection; Radio measurements; Filming crew

Start time: 13:39

End time: 15:28

Narrative: Cesare, Luz Ma and Ben, together with a guest from the filming crew, drove the rover along Cow Dung Rd and Galileo Rd. They took weather observation, radio measurements, and collected rock samples before going to the lower lands along Cactus Rd. Here, they took more measurements before heading back to the hab. An engineering EVA was performed at the end of the excursion

Destination: Galileo / Cactus Rd

Coordinates (use UTM NAD27 CONUS): 12S 520200, 4251500

Operations Report – January 02nd



Crew 218 Operations Report 2-Jan-20

SOL: 11

Name of person filing report: Luz Maria Agudelo Urrego

Non-nominal systems: NA

Generator: run

Hours run: 12.5

From what time last night: 1900

To what time this morning: 0730

List any additional daytime hours when the generator was run: N/A

Solar— SOC % (Before the generator is run at night: 64%

Diesel Reading – 74%

Station Propane Reading – 62%

Ethanol Free Gasoline: N/A

Water (loft tank) (gal): 36

Water Meter (units): 0147236.5

Water (static tank) (gal): 270

Static to Loft Pump used – No

Water in Green Hab (gal): 80

Water in Science Dome (gal): 0

Toilet tank emptied: No

Deimos rover used: Still in the workshop


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: Assigned to director


Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: EVA 10

Hours: 121.7

Beginning charge: (Before EVA): 100%

Ending charge: (On return from EVA, before recharging): 86%

Currently charging: Yes

Spirit rover used: EVA 11

Hours: 122.2

Beginning charge: (Before EVA): 100%

Ending charge: (On return from EVA, before recharging): 57%

Currently charging: Yes

Opportunity rover used: Still in the workshop


Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: EVA 10

Hours: 127

Beginning charge: 100%

Ending charge: (On return from EVA, before charging): 88%

Currently charging: Yes

Curiosity rover used: EVA 11

Hours: 127.5

Beginning charge: 100%

Ending charge: (On return from EVA, before charging): 61%

Currently charging: Yes

Notes on rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use: N/A

Oil Added? No

# Hours the ATVs were used today: 0

Notes on ATVs: N/A

HabCar used, and why, where? No

CrewCar used, and why, where? No

General notes and comments: Nominal

Summary of the internet: Nominal

Summary of suits and radios: We received instruction from Scott Davis about how to fix Suit N°6 on the two-piece space simulator suits. We will fix the suit tomorrow, January 3rd, and let you know how it goes.

Summary of Hab operations: The pipe from the static tank through the loft tank is still frozen. The crew did not fill out the Loft Tank by carrying water buckets from the Static Tank since we still have water in the Loft Tank.

Summary of Science Dome operations: Nominal

Summary of RAM operations: Nominal

Summary of any observatory issues: Nominal

Summary of health and safety issues: Nominal

Questions, concerns, and requests to Mission Support: NA

Luz Ma
Crew 218

Sol Summary – January 02nd


 Sol: 11

Summary Title: Mars Stars

Author's name: Pat Pesa

Mission Status: Can't wait until we recieve our Hollywood Walk of Fame Star!

Sol Activity Summary: Our friendly neighboorhood Martian film crew came to document our daily life and some beautiful Mars landscapes. They followed us on our two EVAs (being very respectable of allowing us to still acomplish our work), and after we shared cookies, mission patches, and stories of travel.

Look Ahead Plan: Tomorrow we will perform our last EVA, and then start cleaning and preparing for the end of our Mission.

Anomalies in work: None

Weather: Partly Cloudy , around 34 degrees

Crew physical status: Healthy

EVA: EVA #10 this morning went to marble ritual for seismic tests. EVA #11 this afternoon went to the same location as EVA 8 of yesterday, Galileo road.

Reports to be filed: sol summary, commander report, operations report, greenhab report, journalist report, EVA report, EVA request, science report

Support Requested: none

Pat Pesa

Geologist, MDRS 218
Copyright © The Mars Society. All rights reserved. | Main Site